The weight, in grams, of the pair of kidneys in adult males between the ages of 40 and 49 are normally distributed with a mean of 327 grams and a standard deviation of 31 grams. For women between 40 and 49, the mean weight of kidneys is 224 g with a standard deviation of 34 grams. A doctor is doing an autopsy on a man and woman who died under suspicious circumstances. The male’s kidneys weighed 256 grams, while the female’s kidneys weighed 299 grams.

Which person’s kidneys were more unusual

Gender

Mean kidney weight

Stand. Dev.

Male

327 grams

31 grams

Female

224 grams

34 grams
Answer #1

z score for male =(X-mean)/standard deviation =(256-327)/31= -2.29

z score for female =(X-mean)/standard deviation =(299-224)/34= 2.21

since absolute value of z score is higher for male , therefore Male kidney is more unusual.